# Learning the Regularization Strength for Deep Fine-Tuning via a Data-Emphasized Variational Objective

### v1

@@ 1 Introduction
Ultimately, we contribute methods that should help practitioners perform cost-effective transfer learning on custom datasets. When available data is plentiful, our experiments suggest our approach is competitive in accuracy while reducing total training time from 16 hours for L2-SP [45] and 150 hours for PTYL [39] (using the grid search ranges recommended by the original authors) to under 3 hours. When available data is limited, e.g., only 5-300 labeled examples per class, our experiments suggest our approach can be particularly effective in improving accuracy and runtime.
@@ 3.1 Variational methods for posterior estimation
$J_{\text{ELBo}}:=\mathbb{E}_{q(w,V)}\left[\log\sum_{i=1}^{N}p(y_{i}|w,V)\right]-\mathbb{KL}(q(w)\|p(w|\lambda))-\mathbb{KL}(q(V)\|p(V|\tau)).$ (4)
@@ 3.3 Inflating dataset size to select complex models with better generalization
$J_{\text{DE-ELBo}}:=\kappa\mathbb{E}_{q(w,V)}\left[\log\sum_{i=1}^{N}p(y_{i}|w,V)\right]-\mathbb{KL}(q(w,V)\|p(w,V))$ (8)
@@
Demonstrating the value of $\kappa=D/N$ for improved selection. In Fig. 1, we compare the cases of $\kappa=1$ (left panel) and $\kappa=D/N$ (right) on CIFAR-10 for L2-SP. In each plot, we compare two possible $q$, “A” (in pink) and “B” (in purple), across a range of $\lambda,\tau$ values. Version $A$ sets $\bar{w}_{A},\bar{V}_{A},\bar{\sigma}_{A}$ to a solution favored by ELBo, with known test-set accuracy 34.9%. Version $B$ sets $\bar{w}_{B},\bar{V}_{B},\bar{\sigma}_{B}$ to values favored by our data-emphasized ELBo, with known test-set accuracy of 87.5%. We see the poor-accuracy version A model is strongly favored by the conventional ELBo, while the more accurate version B model is favored by our recommended DE ELBo with $\kappa=D/N$.
@@ 5 Discussion and Conclusion
When using PTYL, we found that $\mu,\Sigma$ values obtained from supervised pre-training (minimizing cross-entropy on the source task) lead to better transfer learning (better target task accuracy) than using SimCLR [6] self-supervised learning on the source task. Consistent with Špendl and Pirc [40]’s findings on the Oxford Flowers-102 dataset [34] without hyperparameter tuning, we find that Shwartz-Ziv et al. [39]’s supervised prior performs better than their self-supervised prior on CIFAR-10 and Oxford-IIIT Pet (see App. F).

### v2

@@ 1 Introduction
Ultimately, we contribute methods that should help practitioners perform cost-effective transfer learning on custom datasets. When available data is plentiful, our experiments suggest our approach is competitive in accuracy while reducing total training time from 16 hours for L2-SP [45] and 149 hours for PTYL [39] (using the grid search ranges recommended by the original authors) to under 3 hours. When available data is limited, e.g., only 5-300 labeled examples per class, our experiments suggest our approach can be particularly effective in improving accuracy and runtime.
@@ 3.1 Variational methods for posterior estimation
$J_{\text{ELBo}}:=\mathbb{E}_{q(w,V)}\left[\sum_{i=1}^{N}\log p(y_{i}|w,V)\right]-\mathbb{KL}(q(w)\|p(w|\lambda))-\mathbb{KL}(q(V)\|p(V|\tau)).$ (4)
@@ 3.3 Inflating dataset size to select complex models with better generalization
$J_{\text{DE ELBo}}:=\kappa\mathbb{E}_{q(w,V)}\left[\sum_{i=1}^{N}\log p(y_{i}|w,V)\right]-\mathbb{KL}(q(w,V)\|p(w,V))$ (8)
@@
Demonstrating the value of $\kappa=D/N$ for improved selection. In Fig. 1, we compare the cases of $\kappa=1$ (left panel) and $\kappa=D/N$ (right) on CIFAR-10 for L2-SP. In each plot, we compare two possible $q$, “A” (in pink) and “B” (in purple), across a range of $\lambda,\tau$ values. Version $A$ sets $\bar{w}_{A},\bar{V}_{A},\bar{\sigma}_{A}$ to a solution favored by ELBo, with known test-set accuracy 28.6%. Version $B$ sets $\bar{w}_{B},\bar{V}_{B},\bar{\sigma}_{B}$ to values favored by our data-emphasized ELBo, with known test-set accuracy of 87.3%. We see the poor-accuracy version A model is strongly favored by the conventional ELBo, while the more accurate version B model is favored by our recommended DE ELBo with $\kappa=D/N$.
@@ 5 Discussion and Conclusion
When using PTYL, we found that $\mu,\Sigma$ values obtained from supervised pre-training (minimizing cross-entropy on the source task) lead to better transfer learning (better target task accuracy) than using SimCLR [6] self-supervised learning on the source task. Consistent with Špendl and Pirc [40]’s findings on the 102 Category Flower dataset [34] without hyperparameter tuning, we find that Shwartz-Ziv et al. [39]’s supervised prior performs better than their self-supervised prior on CIFAR-10 and Oxford-IIIT Pet (see App. F).